A firm charges $40 per unit for the first three units of a good purchased, and $20 for each additional unit thereafter. The firm’s marginal cost and average total cost are both constant at $15. A consumer purchases six units. How much profit will the firm earn?

Group of answer choices

$70

$110

$90

$80

$50

2. Suppose there are two types of consumers for cell phones and accessories (cases, extra chargers, etc.) Consumers of type A are willing to pay $750 for a phone and $40 for the accessories. Consumers of type B are willing to pay $720 for a phone and $90 for the accessories. The firm selling these products faces no competition and has a marginal cost of zero. What is the optimal commodity bundling strategy?

Group of answer choices

charge $840 for a phone and accessories

charge $790 for a phone and accessories

charge $810 for a phone and accessories

charge $760 for a phone and accessories

3. Which of the following statements is correct regarding the peak-load pricing strategy?

Group of answer choices

The peak price and peak quantity are both higher than the off-peak price and off-peak quantity.

The peak price is lower than the off-peak price, but the peak quantity is higher than the off-peak quantity.

The off-peak price and off-peak quantity are both higher than the peak price and peak quantity.

The peak price is higher than the off-peak price, but the peak quantity is lower than the off-peak quantity.
Answer #1

1) Total revenue = 3*40+ 3*20= 180

Total cost = ATC*Q = 15*6 = 90

Profit = TR-TC = 180-90 = $90

The answer is option (c).